Find the exact solutions and simplify as much as possible: 5x2 − 1 = 2x. Show work.
Answer by Aldorozos(172)

I am assuming you meant 5x^2-2x-1 =0
This is a quadratic equation
X=( 2+ and - sort of 4 +20 )/10
Check the book or Google. Com to learn about quadratic equations.
X= (2+ And minus sqrt of 24)/2 this is the answer however in can be simplified
X= (2+ And minus sqrt of 4*6)/2
X= 2+ And minus (2 sqrt of 6) divided by 10
Factoring out 2,we get (1+and minus sqrt 6)/5